# Navier-Stokes Equations with Fractional Dissipation and Associated Doubly Stochastic Yule Cascades

### v1

@@ Abstract
Parametric regions are identified in terms of the spatial dimension $d$ and the power $\gamma$ of the Laplacian that separate explosive from non-explosive regimes for the self-similar doubly stochastic Yule cascades (DSY) naturally associated with the deterministic fractional Navier-Stokes equations (FNSE) on $\mathbb{R}^{d}$ in the scaling-supercritical setting. Explosion and/or geometric properties of the DSY, are then used to establish non-uniqueness, local existence, and finite-time blow-up results for a scalar partial differential equation associated to the FNSE through a majorization principle at the level of stochastic solution processes. The solution processes themselves are constructed from the DSY and yield solutions to the FNSE upon taking expectations.
In the special case $d=2$, a closed-form expression for the solution process of the FNSE is derived. This representation is employed to prove finite-time loss of integrability of the solution process for sufficiently large initial data. Notably, this lack of integrability does not necessarily imply blow-up of the FNSE solutions themselves. In fact, in the radially symmetric case, solutions can be continued beyond the integrability threshold by employing a modified notion of averaging.
@@ 1 Introduction
One way to approach the unresolved well-posedness issues of the incompressible 3D Navier–Stokes equations is through their natural scaling properties, see (2.1). This perspective highlights, for instance, the existence of a scaling gap between quantities that can be controlled a priori and the functional bounds that are critical or supercritical in nature [bradshaw2019algebraic, lemarie2002recent, tao09, cheskidov2022sharp].
Another manifestation of this scaling issue arises when considering the Navier–Stokes equations on $\mathbb{R}^{d}$, $d\geq 2$, with fractional dissipation $\gamma>0$, given by the Laplacian raised to a fractional power, see $\textup{(NSE)}_{d,\gamma}$. On and above the critical line
@@
there are a number of global-in-time well-posedness and regularity results, including the classical 2D Navier–Stokes equation [lady, constantin_foias, katz, tao09, nan, baaske, ferreira, colombo].
In contrast, below the critical line, that is, in the supercritical regime $\gamma<(d+2)/4$, the questions of global well-posedness and regularity remain largely open, with some evidence suggesting finite-time blow-up and non-uniquesness of solutions [cordoba, DA_EB_MC2022, de2019infinitely]. This supercritical regime includes the physically relevant 3D case with $\gamma=1$. Incidentally, the critical case $\gamma=(d+2)/4$ corresponds precisely to the situation in which the energy, given by the $L^{2}$-norm of the solution, is scaling-invariant; see Section 2, and in particular (2.2).
In this paper we investigate how the interplay between dimension, dissipation, and criticality in the fractional Navier–Stokes problem on $\mathbb{R}^{d}$ can be understood through a naturally associated stochastic structure, first identified by LeJan and Sznitman [lejan]. As shown in Section 2, this structure arises directly from the natural rescaling of the Fourier-transformed fractional Navier–Stokes equations in mild form, see (2.6)–(2.12).
@@
which can be derived by rescaling $\widehat{u_{\lambda}}$ with $\lambda=1/|\xi|$. Thus the exponent $2\gamma-d-1$ is directly linked to the natural scaling (2.1); see [chaos] for detailed calculations in the case $d=3$, $\gamma=1$. Remarkably, the corresponding integral kernel $H$ given by (2.12) is a probability kernel yielding a stochastic interpretation of solutions precisely in the supercritical regime $1/2<\gamma<(d+2)/4$.
In this paper, we look at the connection between of the dimension and dissipation and the nature of criticality of the fractional Navier-Stokes problem in $\mathbb{R}^{d}$ through the prism of naturally associated stochastic structure, first observed by LeJan and Sznitman [lejan]. As we shall see in Section 2, these structures arise directly via natural re-scaling of the Fourier-transformed fractional Navier-Stokes equations in a mild formulation, see (2.6)–(2.12). It is important to note that the scaling kernel $h(\xi)=c_{d,\gamma}|\xi|^{2\gamma-d-1}$ can be obtained via rescaling $\widehat{u_{\lambda}}$ with $\lambda=1/|\xi|$. Thus the exponent $2\gamma-d-1$ has a direct connection to the natural scaling (2.1), see [chaos] for a detailed calculations in the case $d=3$, $\gamma=1$. Moreover, as it turns out, the resulting integral kernel $H$ is a probability kernel yielding the stochastic interpretation of solutions precisely in the supercritical regime $1/2<\gamma<(d+2)/4$.
Indeed, the resulting equation (nFNS) can be interpreted as a mean flow equation associated with a stochastic structure that can be described in terms of two random fields:
@@
(i) a doubly-stochastic Yule (DSY) random field $\mathbb{Y}$, indexed by a labelled binary tree. The tree-indexed labels form a Markov process when restricted to any fixed ray emanating from the root, as discussed in Section 3; and
@@
(ii) a stochastic solution process $\bf{X}$}, defined via a binary recursion on the evolution of $\mathbb{Y}$ within a common probability space $(\Omega,\mathscr{F},P)$, see Section 2.
@@
The expectation of $\bf{X}$} then satisfies the associated mean field equation, which provides the framework for studying various regularity questions for the mean flow, such as uniqueness versus non-uniqueness, finite-time blow-up, and related phenomena.
@@
This framework leads to a natural unified approach to the incompressible fractional Navier–Stokes equations on $\mathbb{R}^{d}$, within the scaling-invariant functional setting $u\in L^{\infty}([0,T),PM^{a}),$ where $PM^{a}$ denotes the space of pseudomeasures defined in (2.3), with $a=d+1-2\gamma$.
@@
Well-posedness the three-dimensional Navier–Stokes equations in such settings was first studied in [lejan] using probabilistic approach, and later in [lemarie2002recent] via analytical methods, establishing global existence of mild solutions with small initial data. More recently, [ferreira] proved a number of local and global regularity results for mild solutions with initial data $u_{0}\in PM^{a,b}:=PM^{a}+PM^{b}$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$.
@@
As mentioned above, our goal is to focus on phase transitions of the stochastic structures $\bf{X}$} and $\mathbb{Y}$, and to explore their possible connections to regularity problems; see, for example, [BoutrosGibbon, edengrowth, chaos]. Several key elements of the analysis include: (a) the stochastic geometry underlying the representation of the solution process $\bf{X}$}; (b) stochastic symmetries of both $\mathbb{Y}$ and $\bf{X}$} under groups of tree automorphisms; (c) critical parameters governing stochastic explosion events for $\mathbb{Y}$; and (d) integrability of the (minimal) solution process, and exploiting symmetries to continue solutions beyond the existence of mean flow.
@@
On the one hand, we identify parametric regions in terms of $d$ and $\gamma$ where the DSY cascade $\mathbb{Y}$ exhibits either explosive or non-explosive behavior; see Section 3 and Figure 4. For the fractional Navier–Stokes equation itself, stochastic explosion alone does not qualitatively strengthen existing existence and uniqueness results for mild solutions (see Section 4 and the subsequent remark). However, in the scalar case of the Montgomery–Smith equation introduced in [smith] and its fractional version (MS)-(FMS) – the geometric structure of $\bf{X}$} prevents cancellations, so that explosive regimes naturally lead to non-uniqueness of solutions; see Section 4.
Interestingly, in both the explosive and non-explosive cases, the mean flows $\mathbb{E}({\bf{X}})$ exhibit finite-time blow-up [dascaliuc2019jan]. Although the arguments in [dascaliuc2019jan] concerned a different kernel and the specific case $d=3$, $\gamma=1$, they can be readily adapted to the fractional Montgomery–Smith equation on $\mathbb{R}^{d}$ across the entire supercritical range $\frac{1}{2}<\gamma<\frac{d+2}{4}.$
@@
Perhaps counterintuitively, the explosive regime for $\mathbb{Y}$ occurs near the critical line $\gamma=(d+2)/4$ (see Figure 4), where existence and uniqueness results are more plausible and partial regularity theory is available (see, e.g., [colombo]). By contrast, the non-explosive regimes lie deeper in the supercritical region, where well-posedness results are expected to be weaker. This highlights the limitations of interpreting stochastic explosion of the DSY cascade as a direct metaphor for ill-posedness.
@@
On the other hand, in the two-dimensional case we exploit the geometric structure of the bilinear term in the Navier–Stokes nonlinearity to obtain a closed-form expression for the minimal solution process $\bf{X}$} (see Section 5). This allows us to establish symmetry conditions on the initial data which, together with the symmetries of the DSY cascade $\mathbb{Y}$, make it possible use a modified notion of averaging to extend solutions beyond the critical time $t_{c}$ that marks the point at which $\bf{X}$} ceases to be integrable; see Section 5. We further show that in 2D, the critical time $t_{c}$ is finite when the initial data are sufficiently large (see (5)).
@@
While it is classical that rotationally invariant initial data in 2D lead to global solutions via cancellations in the nonlinearity, our result is the first to reveal this cancellation directly at the level of the stochastic cascades. This perspective allows us to define the “mean evolution” of $\bf{X}$} beyond the integrability threshold $t_{c}$, thereby opening a pathway to bypass the limitations imposed by the use of majorizing principles (see Section 4), which have played a central role in earlier existence proofs [lejan, dascaliuc2019jan].
@@
Nevertheless, a key challenge remains: the discrepancy between the loss of integrability of $\bf{X}$} (expressed by the finiteness of $t_{c}$) and the blow-up of the mean-flow equation. Addressing this gap requires developing a more general framework for continuing the evolution past $t_{c}$, perhaps by introducing new notions of averaging for $\bf{X}$} that exploit cancellations arising from symmetries in both the solution process $\bf{X}$} and the DSY cascade $\mathbb{Y}$.
@@
The paper is organized as follows. In Section 2 the incompressible fractional Navier-Stokes equations are introduced, both as a PDE and in terms of the stochastic structures: the DSY $\mathbb{Y}$ and the (minimal) solution process $\bf{X}$}. Their relationship via the mean flow equation is established in Proposition 2, where the notion of a minimal solution process and integrable solution process naturally appear. The latter notion of minimality involves a notion of stochastic explosion that is defined here as well. The general concept of solution process stands apart from that of an integrable solution process and is defined via a binary stochastic recursion that transforms initial data in accordance with the ‘bilinear product’ intrinsic to the incompressible mean flow. The stochastic/geometric structure of the DSY $\mathbb{Y}$ is analyzed in Section 3. In particular, stochastic explosion/non-explosion events are characterized in terms of the dimension $d$ and the fractional exponent parameters $\gamma$. In Section 4 the Montgomery-Smith model is introduced as a tool for obtaining estimates on the critical time governing integrability of the solution process via a comparison method, more generally developed in [dascaliuc2019jan] for the incompressible Navier-Stokes equation. Section 5 involves a further specialization to $d=2$ dimensions, where the stochastic geometry and symmetries underlying the solution process $\bf{X}$} are significantly simplified, while preserving the key elements and critical phenomena noted above. The role of stochastic ‘cancellations’ due to symmetries in the initial data and stochastic solution process are a special focus of this paper, and are shown to lead simply to a closed form solution defined over the critical range of integrability. Interestingly, it is also shown that this locally defined solution extends to a global solution.
@@ 2 Navier-Stokes equations with Fractional Dissipation
$u_{\lambda}(x,t)=\lambda^{2\gamma-1}u(\lambda x,\lambda^{2\gamma}t),\ \ p_{\lambda}(x,t)=\lambda^{4\gamma-2}p(\lambda x,\lambda^{2\gamma}t),\ \ u_{0\lambda}(x)=\lambda^{2\gamma-1}u(\lambda x).\ \$(2.1)
@@
is the energy. Under the natural scaling (2.1), the energy is scaled as
$E[u_{\lambda}](t)=\lambda^{4\gamma-(d+2)}E[u](\lambda^{2\gamma}t).$
The system $\textup{(NSE)}_{d,\gamma}$ is called critical if $\gamma=\frac{d+2}{4}$ (i.e. if the energy is scale-invariant), supercritical if $\gamma<\frac{d+2}{4}$, and subcritical if $\gamma>\frac{d+2}{4}$. The system is called hypodissipative for $\gamma<1$ and hyperdissipative for $\gamma>1$. Global regularity is obtained for the critical and subcritical regime $\gamma\geq\frac{d+2}{4}$ (see e.g. [katz, tao09, nan]). Local and global regularity holds for mild solutions for any $\gamma\in(\frac{1}{2},\frac{d+2}{4})$ and $u_{0}\in PM^{a,b}:=PM^{a}+PM^{b}$ where $PM^{a}$ is the space of pseudomeasure-type [ferreira]
@@
Local and global regularity of mild solutions in the hyperdissipative regime $\gamma\geq 1$ for initial data $u_{0}$ in supercritical Besov spaces and Triebel-Lizorkin spaces are obtained in [baaske].
The case of particular interest is $d=3$, in which partial regularity is obtained for the supercritical hyperdissipative regime $1<\gamma<5/4$ [katz]. If $u_{0}\in H^{\delta}$ for some arbitrary $\delta>0$, then global regularity holds for any $\gamma\in[\frac{5}{4}-\epsilon,\frac{5}{4}]$ where $\epsilon>0$ depends on $\|u_{0}\|_{H^{\delta}}$ [colombo]. Finite-time blowup of solutions to a forced system is obtained in [cordoba] for the hypodissipative regime $0\leq\gamma<\frac{2}{9}(22-8\sqrt{7})$.
The PDE $\textup{(NSE)}_{d,\gamma}$ has an integral form
@@
Taking the Fourier transform of (2.4) and using the identity $\mathscr{F}\{fg\}=c_{0}\mathscr{F}\{f\}*\mathscr{F}\{g\}$ (see [folland, p. 284]), one has
@@
where $e_{\xi}=\xi/|\xi|$ is the unit vector in $\xi$-direction and $\pi_{\xi^{\perp}}$ is the perpendicular projection onto the hyperplane orthogonal to $\xi$. As remarked in [dascaliuc2019jan], the vector product $\otimes$ in (nFNS) can be replaced by a symmetrized vector product $\odot$ defined by
@@
For $\gamma\in(\frac{1}{2},\frac{d+2}{4})$, one can infer from the Fourier transform of radial functions [rudin, p. 205]
@@
The function $h$ defined by (2.10) is known as a standard majorizing kernel [rabi]. The normalized Fourier transform $\chi(\xi,t)=c_{0}\frac{\hat{u}(\xi,t)}{h(\xi)}$ satisfies
@@
At this point, the solution $\chi$ to (nFNS) can be interpreted probabilistically as the expected value of a “solution process” $\mathbf{X}$ defined implicitly by
@@
We note that (2.14) can also be viewed as a recursion defined on a Doubly Stochastic Yule Cascade (DSY): $\mathbb{Y}=\{Y_{v}\}_{v\in\mathbb{T}}$ built on a binary tree $\mathbb{T}=\{\theta\}\bigcup\left(\cup_{n=1}^{\infty}\{1,2\}^{n}\right)$ rooted at $\theta$, where $Y_{\theta}=Y$ is an exponential random variable with intensity $|\xi|^{2}$, $Y_{1}$ and $Y_{2}$ are the exponential random variables with intensities $|W_{1}|^{2}$ and $|W_{2}|^{2}$, and so on down the tree $\mathbb{T}$, see Section 3 and Section 3 for a more complete definition. In this context, the following two parameters, one stochastic and the other deterministic, play an essential role in construction of a well-defined solution process.
@@
The (stochastic) explosion time is defined by
@@
where for each vertex $v\in\mathbb{T}$, we denote by $|v|$ the genealogical height of vertex $v$, and $v|j$ denotes the truncation up to the $j$’th generation with the convention that $v|0=\theta$.
The critical parameter
@@
The no-explosion event $S=\infty$ a.s. implies that for any $t\geq 0$, the random tree implicit to the iteration of the recursion (2.14) is a.s. finite. In particular, since any solution process can be factored as an ${\otimes}_{\xi}$}-product, the minimality follows by induction. ∎
@@
The proof is very similar to [dascaliuc2019jan, Prop. 2.2]. Since $\mathbf{X}(\xi,t)\in L^{1}(\mathbb{P}_{\xi})$, the function $\chi(\xi,t)=\int_{\Omega}\mathbf{X}(\xi,t,\omega)\mathbb{P}_{\xi}(d\omega)$ is measurable on $Q$ (see [folland, Thm 2.39]). To show (2.17), we take the expected value of the magnitude of both side of (2.14):
@@
Based at the origin, vector $\xi$ partitions the plane into two halves, labeled as and in Figure 1.

### v2

@@ Abstract
We identify a probabilistic structure known as self-similar doubly stochastic Yule (DSY) cascade associated with the deterministic Navier-Stokes equations (NSE) in $\mathbb{R}^{d}$ with fractional dissipation $(-\Delta)^{\gamma}$. Interestingly, such a structure is well-defined only in the scaling-supercritical regime $\gamma\in(\frac{1}{2},\frac{d+2}{4})$. We then characterize parametric regions of $(d,\gamma)$ that correspond to the non-explosive, explosive, and hyperexplosive behavior of the DSY cascade. Stochastic solution processes are constructed recursively, and their expectations yield solutions to the fractional NSE whenever these expectations exist. Explosion and geometric properties of the DSY cascade are then exploited to establish non-uniqueness and finite-time blowup results for a scalar partial differential equation associated with the fractional NSE thanks to a majorization principle for stochastic solution processes. In the special case $d=2$, we derive a closed form for the solution process and prove the finite-time loss of integrability of the solution process for sufficiently large initial data. Notably, this lack of integrability does not necessarily imply finite-time blowup of solutions to the fractional NSE. Indeed, for vortex-flow initial data, we show that the solution can be continued beyond the time of integrability breakdown by employing a notion of averaging.
@@ 1 Introduction
One way to approach the unresolved well-posedness issues of the incompressible 3D Navier–Stokes equations is through their natural scaling properties; see (2.1). This perspective highlights, for example, the existence of a scaling gap between quantities that can be controlled a-priori and the functional bounds that are critical or supercritical in nature; [8, 29, 38, 10].
Another manifestation of this scaling issue arises when considering the Navier–Stokes equations on $\mathbb{R}^{d}$, $d\geq 2$, with fractional dissipation $\gamma>0$, given by the Laplacian raised to a fractional power; see $\textup{(NSE)}_{d,\gamma}$. On and above the critical line
@@
there are a number of global-in-time well-posedness and regularity results, including the classical 2D Navier–Stokes equation; [27, 12, 26, 38, 33, 3, 23, 11].
In contrast, below the critical line, that is, in the supercritical regime $\gamma<(d+2)/4$, the questions of global well-posedness and regularity remain largely open, with some evidence suggesting finite-time blow-up and non-uniqueness of solutions [13, 1, 21]. This supercritical regime includes the physically relevant 3D case with $\gamma=1$. Incidentally, the critical case $\gamma=(d+2)/4$ corresponds precisely to the situation in which the energy, given by the $L^{2}$-norm of the solution, is scaling-invariant; see Section 2, and in particular (2.2).
While the most profound example relating partial differential equations and stochastic processes in mathematical physics is no doubt that of the heat/diffusion equation and Brownian motion, other somewhat surprising illustrations are known in other contexts, including KPP and wave equations [31, 9, 4], discounted Riccati-type equations [2, 20, 17, 19], and incompressible Navier-Stokes equations [28]. In this paper we investigate how the interplay between dimension, dissipation, and criticality in the fractional Navier–Stokes problem on $\mathbb{R}^{d}$ can be understood through a naturally associated stochastic structure originating with LeJan and Sznitman [28], where a scaling kernel $h(\xi)$ is naturally associated with the Fourier transformed equation and wave number $\xi$. As shown in Section 2, this structure arises directly from the natural rescaling of the Fourier-transformed fractional Navier–Stokes equations in the mild form; see (2.6)–(2.12).
@@
which can be derived by rescaling $\widehat{u_{\lambda}}$ with $\lambda=1/|\xi|$. Thus, the exponent $2\gamma-d-1$ is directly linked to the natural scaling (2.1); see [14] for detailed calculations in the case $d=3$, $\gamma=1$. Remarkably, the corresponding integral kernel $H$ given by (2.12) is a probability kernel yielding a stochastic representation of solutions precisely in the supercritical regime $1/2<\gamma<(d+2)/4$; here ”criticality” refers to the scaling invariance of the associated energy functional (2.2). Indeed, in this regime, the resulting equation (nFNS) can be interpreted as a mean flow equation associated with a stochastic structure that can be described in terms of two random fields:
@@
(i) A doubly-stochastic Yule (DSY) random field $\mathbb{Y}$, indexed by the vertices of a labeled binary tree. The tree-indexed labels form a Markov process when restricted to any fixed ray emanating from the root, as discussed in Section 3, and
@@
(ii) a stochastic solution process $\bf{X}$}, defined via a binary recursion on the evolution of $\mathbb{Y}$ on a common probability space $(\Omega,\mathscr{F},P)$; see Section 2.
@@
The expected value of $\bf{X}$}, when it exists, satisfies the associated mean field equation, which provides the framework for studying various regularity questions for the mean flow, such as uniqueness versus non-uniqueness, finite-time blow-up, and related phenomena.
@@
This framework leads to a natural unified approach to the incompressible fractional Navier–Stokes equations in $\mathbb{R}^{d}$, within the scaling-invariant functional setting $u\in L^{\infty}([0,T),PM^{a}),$ where $PM^{a}$ denotes the space of pseudo-measures defined in (2.3), with $a=d+1-2\gamma$.
@@
Well-posedness of the three-dimensional Navier–Stokes equations in such settings was first studied in [28] using a closely related probabilistic approach, and later in [29] via analytical methods, establishing global existence of mild solutions with small initial data. More recently, [23] proved a number of local and global regularity results for mild solutions with initial data $u_{0}\in PM^{a,b}:=PM^{a}+PM^{b}$ and $\gamma\in(\frac{1}{2},\frac{d+2}{4})$.
In this paper, we look at the connection between the dimension and dissipation and the nature of criticality of the fractional Navier-Stokes problem in $\mathbb{R}^{d}$ through the prism of this naturally associated stochastic structure. In the original approach, [28] introduced a coin-tossing device to eliminate the possibility of there being infinitely many branchings of the cascade in finite time, to be referred to as (stochastic) explosion. This coin tossing device was removed in [14], opening the possibility of using stochastic explosion as a mechanism for establishing non-uniqueness of solutions; [20, 19, 18].
@@
Our goal is to focus on phase transitions of the stochastic structures $\bf{X}$} and $\mathbb{Y}$ depending on parameters $d$ and $\gamma$, and to explore their possible connections to regularity problems; see, for example, [7, 22, 14]. Several key elements of the analysis include: (a) The stochastic geometry underlying the representation of the solution process $\bf{X}$}; (b) Stochastic symmetries of both $\mathbb{Y}$ and $\bf{X}$} under groups of tree automorphisms; (c) Critical parameters governing stochastic explosion and hyperexplosion events for $\mathbb{Y}$, see Definitions 2 and 2; and (d) Integrability of the (minimal) solution process, while exploiting symmetries to continue solutions beyond the existence of the mean flow.
@@
First, we identify parametric regions in terms of $d$ and $\gamma$ where the DSY cascade $\mathbb{Y}$ exhibits either explosive or non-explosive behavior; see Section 3.2 and Figure 6. The numerical evidence shows existence of hyperexplosive regions for any dimension $d$, based on the new criteria, see Section 3.2, that hold for general self-similar DSY cascades as defined in [16]. This opens the possibility in the future to expand the probabilistic approach developed for the $\alpha$-Riccati equation, [20, 19], to the fractional Navier-Stokes equations. The former relies heavily on stochastic hyperexplosion, but for the fractional Navier–Stokes equation itself, stochastic explosion/hyperexplosion alone does not qualitatively strengthen existing existence and uniqueness results for mild solutions; see Section 4 and the subsequent remark. However, in the scalar case of the Montgomery–Smith equation, introduced in [32], and its fractional version (MS)-(FMS), the geometric structure of $\bf{X}$} prevents cancellations, so that explosive regimes naturally lead to non-uniqueness of solutions; see Section 4. Interestingly, in both the explosive and non-explosive cases, the mean flows $\mathbb{E}({\bf{X}})$ exhibit finite-time blow-up; [18]. Although the arguments in [18] concern a different kernel, in the specific case $d=3$, $\gamma=1$, they can be readily adapted to the fractional Montgomery–Smith equation on $\mathbb{R}^{d}$ across the entire supercritical range $\frac{1}{2}<\gamma<\frac{d+2}{4}.$
@@
Perhaps counterintuitively, the explosive and especially hyperexplosive regimes for $\mathbb{Y}$ occur near the critical line $\gamma=(d+2)/4$ (see Figure 6), where existence and uniqueness results are more plausible and partial regularity theory is available (see, e.g., [11]). By contrast, the non-explosive regimes lie deeper in the supercritical region, where well-posedness results are expected to be weaker. This highlights the limitations of interpreting stochastic explosion of the DSY cascade as a direct metaphor for ill-posedness.
@@
Another main takeaway of this research is that in the two-dimensional case the geometric structure of the bilinear term in the Navier–Stokes nonlinearity can be exploited to obtain a closed-form expression for the minimal solution process $\bf{X}$} (see Section 5). This permits one to establish symmetry conditions on the initial data which, together with the symmetries of the DSY cascade $\mathbb{Y}$, make it possible to use a modified notion of averaging to extend solutions beyond the critical time $t_{c}$ that marks the point at which $\bf{X}$} ceases to be integrable; see Section 5. We further show that in the case $d=2$, the critical time $t_{c}$ is finite when the initial data are sufficiently large; see (5).
@@
While it is classical that rotationally invariant initial data in 2D lead to global solutions via cancellations in the nonlinearity, the result presented here is the first to reveal this cancellation directly at the level of the stochastic cascades. This perspective allows us to define the “mean evolution” of $\bf{X}$} beyond the integrability threshold $t_{c}$, thereby opening a pathway to bypass the limitations imposed by the use of majorizing principles (see Section 4), which have played a central role in earlier existence proofs [28, 18].
@@
Nevertheless, a key challenge remains: There is a discrepancy between the loss of integrability of $\bf{X}$} (expressed by the finiteness of $t_{c}$) and the blow-up of the mean-flow equation. To address this gap requires development of a more general framework for continuing the evolution past $t_{c}$, perhaps by introducing new notions of averaging for $\bf{X}$} that exploit cancellations arising from symmetries in both the solution process $\bf{X}$} and the DSY cascade $\mathbb{Y}$. An illustrative result along these lines is provided in Section 5.
@@
The paper is organized as follows. In Section 2 the incompressible fractional Navier-Stokes equations are introduced, both as a PDE and in terms of the stochastic structures: the DSY $\mathbb{Y}$ and the (minimal) solution process $\bf{X}$}. Their relationship via the mean flow equation is established in Section 2, where the notion of a minimal solution process and integrable solution process naturally appear. The latter notion of minimality involves a notion of stochastic explosion that is defined here as well. The general concept of solution process stands apart from that of an integrable solution process and is defined via a binary stochastic recursion that transforms initial data in accordance with the ‘bilinear product’ intrinsic to the incompressible mean flow. The stochastic/geometric structure of the DSY $\mathbb{Y}$ is analyzed in Section 3. In particular, stochastic explosion/non-explosion events are characterized in terms of the dimension $d$ and the fractional exponent parameters $\gamma$. In Section 4 the Montgomery-Smith model is introduced as a tool for obtaining estimates on the critical time governing integrability of the solution process via a comparison method, more generally developed in [18] for the incompressible Navier-Stokes equation. Section 5 involves a further specialization to $d=2$ dimensions, where the stochastic geometry and symmetries underlying the solution process $\bf{X}$} are significantly simplified, while preserving the key elements and critical phenomena noted above. The role of stochastic ‘cancellations’ due to symmetries in the initial data and stochastic solution process are a special focus of this paper, and are shown to lead simply to a closed form solution defined over the critical range of integrability. Interestingly, it is also shown that this locally defined solution extends to a global solution.
@@ 2 Navier-Stokes equations with Fractional Dissipation
$u_{\lambda}(x,t)=\lambda^{2\gamma-1}u(\lambda x,\lambda^{2\gamma}t),\ \ p_{\lambda}(x,t)=\lambda^{4\gamma-2}p(\lambda x,\lambda^{2\gamma}t),\ \ u_{0\lambda}(x)=\lambda^{2\gamma-1}u_{0}(\lambda x).\ \$(2.1)
@@
is the energy. Under the natural scaling (2.1), the energy is scaled as $E[u_{\lambda}](t)=\lambda^{4\gamma-(d+2)}E[u](\lambda^{2\gamma}t)$. The system (NSE)d,γ is called critical if $\gamma=\frac{d+2}{4}$ (i.e. if the energy is scale-invariant), supercritical if $\gamma<\frac{d+2}{4}$, and subcritical if $\gamma>\frac{d+2}{4}$. The system is called hypodissipative for $\gamma<1$ and hyperdissipative for $\gamma>1$. Global regularity is known for the critical and subcritical regime $\gamma\geq\frac{d+2}{4}$ (see e.g. [26, 38, 33]). Local and global regularity holds for mild solutions for any $\gamma\in(\frac{1}{2},\frac{d+2}{4})$ and $u_{0}\in PM^{a,b}:=PM^{a}+PM^{b}$ where $PM^{a}$ is the space of pseudomeasure-type [23]
@@
Local and global regularity of mild solutions in the hyperdissipative regime $\gamma\geq 1$ for initial data $u_{0}$ in supercritical Besov spaces and Triebel-Lizorkin spaces are obtained in [3].
The case of particular interest is $d=3$, in which partial regularity is known for the supercritical hyperdissipative regime $1<\gamma<5/4$ [26]. If $u_{0}\in H^{\delta}$ for some arbitrary $\delta>0$, then global regularity holds for any $\gamma\in[\frac{5}{4}-\epsilon,\frac{5}{4}]$ where $\epsilon>0$ depends on $\|u_{0}\|_{H^{\delta}}$ [11]. Finite-time blowup of solutions to a forced system is obtained in [13] for the hypodissipative regime $0\leq\gamma<\frac{2}{9}(22-8\sqrt{7})$.
$\textup{(NSE)}_{d,\gamma}$ has an integral form
@@
Taking the Fourier transform of (2.4) and using the convolution identity $\mathscr{F}\{fg\}=c_{0}\mathscr{F}\{f\}*\mathscr{F}\{g\}$ (see [24]*p. 284), one has
@@
where $e_{\xi}=\xi/|\xi|$ is the unit vector in $\xi$-direction and $\pi_{\xi^{\perp}}$ is the perpendicular projection onto the hyperplane orthogonal to $\xi$. As remarked in [18], the vector product $\otimes$ in (nFNS) can be replaced by a symmetrized vector product $\odot$ defined by
@@
For $\gamma\in(\frac{1}{2},\frac{d+2}{4})$, one can infer from the Fourier transform of radial functions [36, p. 205]
@@
The function $h$ defined by (2.10) is known as a standard majorizing kernel [6]. The normalized Fourier transform $\chi(\xi,t)=c_{0}\frac{\hat{u}(\xi,t)}{h(\xi)}$ satisfies
@@
At this point, the solution $\chi$ to (nFNS) can be interpreted probabilistically as the expected value of a “solution process” $\mathbf{X}$ defined implicitly by the recursive relation
@@
We note that (2.14) can also be viewed as a recursion defined on a Doubly Stochastic Yule Cascade (DSY): $\mathbb{Y}=\{Y_{v}\}_{v\in\mathbb{T}}$ indexed by a binary tree $\mathbb{T}=\{\theta\}\bigcup\left(\cup_{n=1}^{\infty}\{1,2\}^{n}\right)$ rooted at $\theta$, where $Y_{\emptyset}=Y$ is an exponential random variable with intensity $|\xi|^{2}$, $Y_{1}$ and $Y_{2}$ are, conditionally on $Y_{\emptyset}$, the exponential random variables with intensities $|W_{1}|^{2}$ and $|W_{2}|^{2}$, and so on down the tree $\mathbb{T}$, see Section 3 and Section 3 for a more complete definition. In this context, the following two parameters, one stochastic and the other deterministic, play an essential role in the construction of a well-defined solution process.
@@
(i) The (stochastic) explosion time is defined by
@@
where for each vertex $v\in\mathbb{T}$, we denote by $|v|$ the genealogical height of vertex $v$, and $v|j$ denotes the truncation up to the $j$’th generation with the convention that $v|0=\theta$. (ii) The critical parameter given by
@@
The no-explosion event $S=\infty$ a.s. implies that for any $t\geq 0$, the random tree implicit to the iteration of the recursion (2.14) is a.s. finite. In particular, since any solution process can be factored as an ${\odot}_{\xi}$}-product, the minimality follows by induction. ∎
Remark 2.5.
In the explosive cases (i.e. $S<\infty$ a.s.), non-minimal solution processes may exist, particularly in the case of scalar mean flow equations with multiplicative non-linearity, [20, 19]. Crucially, in such cases the following notion of hyper-explosion can play an essential role.
Definition 2.6.
The (stochastic) hyper-explosion time is defined by
$L=L_{\xi}=\underset{n\geq 0}{\mathop{\sup}}\,\underset{|v|=n}{\mathop{\max}}\,\sum\limits_{j=0}^{n}Y_{v|j},$ (2.17)
A DSY cascade is called hyper-explosive if $\mathbb{P}[L<\infty]>0$, and the event $L<\infty]$ is called hyper-explosion.
Remark 2.7.
Clearly, hyper-explosive cascades are explosive. Moreover, by the arguments involving the evolution of $\mathbb{P}(L<t)$ and $\mathbb{P}(S>t)$, see e.g. [14], one can show that in the case of DSY described in Section 3, both explosion and hyper-explosion are zero-one events for any initial state $\xi$.
@@
The proof is very similar to [18, Prop. 2.2]. Since $\mathbf{X}(\xi,t)\in L^{1}(\mathbb{P}_{\xi})$, the function $\chi(\xi,t)=\int_{\Omega}\mathbf{X}(\xi,t,\omega)\mathbb{P}_{\xi}(d\omega)$ is measurable on $Q$ (see [24, Thm 2.39]). To show (2.18), we take the expected value of the magnitude of both side of (2.14):
@@
Based at the origin, the vector $\xi$ partitions the plane into two halves, labeled as and in Figure 1.